OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. Which recursive equation represents the pattern? an = an – 1 + 3n – 2 an = an – 1 + 2n – 3 an = an – 1 + (3n – 2) an = an – 1 + (2n – 3)

OpenStudy (anonymous):

a_n = a_n – 1 + 3^n – 2 a_n = a_n – 1 + 2^n – 3 a_n = a_n – 1 + (3n – 2) a_n = a_n – 1 + (2n – 3)

OpenStudy (anonymous):

Nvm I found it, it was D

ganeshie8 (ganeshie8):

Yep!

ganeshie8 (ganeshie8):

\(\large a_{n-1} = (n-2)^2\) \(\large \implies a_{n} = (n-1)^2 = a_{n-1}-(-2n+3)\)